OpenStudy (anonymous):

THE CIRCUMFERENCE OF A CIRCLE IS 100 INCHES. FIND THE LENGTH OF A RADIUS. OF A CIRCLE. PIR=3.14

OpenStudy (anonymous):

circumference=2.pi.r = so 100=2*3.14*r

OpenStudy (anonymous):

15

OpenStudy (phi):

can you do 2*3.14 ?

OpenStudy (anonymous):

aproxiamtely @nia_maepenix

OpenStudy (anonymous):

6.28?

OpenStudy (phi):

so replace 2*3.14 with 6.28 in your equation. what do you get?

OpenStudy (anonymous):

100/6.28?

OpenStudy (anonymous):

The choices are 63.7 31.8 31.5 10.0

OpenStudy (phi):

your equation becomes 100= 6.28*r as you showed, divide both sides by 6.28, so your equation is 100/6.28 = r you can use a calculator, or type 100/6.28= in google's window r= 15.92 so if that is not one of your choices (roughly) then this is a bogus question notice that the diameter (2 * r) is bout 31.8 so maybe they want the diameter? or a typo in the answer choices.